OpenStudy (anonymous):

what is the degree for this polynomial:

OpenStudy (anonymous):

\[x^3-5x+4\sqrt{x}\]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

how is it 3?

OpenStudy (anonymous):

when looking for a degree of a polynomial you look for the largest exponent

OpenStudy (anonymous):

in your polynomial

OpenStudy (anonymous):

what does the \[4\sqrt{x}\] stand for

OpenStudy (anonymous):

Is it a polynomial at all??? the variable has a fraction power too

OpenStudy (anonymous):

your polynomial is x^3-5x^1+4x^(1/2)

OpenStudy (anonymous):

yes it is a polynomial, im getting this out of my textbook.

OpenStudy (anonymous):

what does the \[4\sqrt{x}\] stand for though

OpenStudy (anonymous):

check again as a polynomial never has fraction or negative power for the variable and in yr question x has a root ie 1/2 power

OpenStudy (anonymous):

So your exponents are {3,1,1/2}. Largest of which are 3; that your degree

OpenStudy (anonymous):

oh ok...thanks

OpenStudy (anonymous):

@imran How can this b a polynom when variable has fraction power??

OpenStudy (anonymous):

I have it written in my book lol

OpenStudy (anonymous):

sanasaj if u look up the definition of polynomial then u will find that either this question is wrong or u hv typed the root wrong

OpenStudy (anonymous):

no it says : \[x^3-5x+4\sqrt{x}\]

OpenStudy (anonymous):

It is not a polynomial in that sense but I think they are just practicing to find degree

OpenStudy (anonymous):

Then pls check with yr teacher as this is not a polynomial but as imran says if it is just for practice then degree is 3

myininaya (myininaya):

by the way its not a polynomial

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

if ur question is answered satisfactorily, pls click on the medal button so that it is shown as a closed question